What would be the 14C activity in a 2600−year old wooden object? Assume that the 14C activity of an equal mass of fresh-cut wood is 13.9 disintegration per second.
1
Expert's answer
2020-04-27T01:52:16-0400

"N\/No = e^(-0.693 T\/Tk)"

N/14=e(0.693*2600/13.9)

N=(1.96*1056 )*14=2.75*1057